OpenStudy (anonymous):

15) Write an equation that is parallel to y = -2x + 1 and passes through the point (1,5)

OpenStudy (anonymous):

Can you identify the slope of y=-2x+1

OpenStudy (anonymous):

For parallel lines, the slopes are the same. For the first equation, the slope is -2, so for the second one, it needs to be the same. You need to use point slope form/equation to figure out the equation. Point Slope: y – y1 = m(x – x1) (m is slope) Plug in the points given: y-5=2(x-1) y-5= 2x-2 y=2x+3 <-- final answer Did this help?

OpenStudy (anonymous):

agree ^^^

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

That's not what I got...

OpenStudy (anonymous):

How did you do it?

OpenStudy (anonymous):

The same as you, just check your 'm' You put 2 instead of -2

OpenStudy (anonymous):

My bad. It would then be y= -2x+7, right?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

Okay, thanks